OpenStudy (anonymous):

Write an equation of a circle with center (4,-7) and r=10

OpenStudy (anonymous):

go right to the answer \[(x-h)^2+(y-k)^2=r^2\] \[(x-4)^2+(y+7)^2=100\]

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

is that the answer